the product table to the right shows that there are thirty-six possible results when two ordinary dice are rolled; the first one is red, and the second is green. Determine the number of possible results for which the sum for both dice is between 2 and 6
Answer #1

number of possible results for which the sum for both dice is between 2 and 6

=N(in which sum is 2)+N(in which sum is 3)+N(in which sum is 4)+N(in which sum is 5)+N(in which sum is 6)=1+2+3+4+5=15

(Note: if 2 and 6 are exclusive ; then number of possible results are 2+3+4 =9)